An arrow is shot at an angle of 85º from the horizontal with an initial velocity of 15m/s.


a.What is its velocity after 9.5 s?



1
Expert's answer
2021-10-31T12:41:27-0400

V=U2-2gh

V= (15sin80)2-2×9.5×h

V=182.25-19h

0=182.25-19h

182.25=19h

=9.61m/s